OpenStudy (anonymous):

Please Help ! The population of Parellel double every 30 years. Its current population is 50,000. Write a mathematical model for the population as a function of time from now, measured in 30-year periods

OpenStudy (wolf1728):

Population = 50,000 *(2^(y/30)) So after 30 years, population = 50,000 *2^(30/30) = 50,000 * 2^1 = 100,000

OpenStudy (anonymous):

thanks how can I rewrite it so that time is measured in years ?

OpenStudy (wolf1728):

Time is measured in years. example after 15 years population equals 50,000 *(2^(y/30)) = 50,000 *(2^(15/30)) = 50,000 *(2^(.5)) = 50,000 *(1.4142135624) =70,711.

OpenStudy (anonymous):

Yeah but the first equation is measured in the 30 year increments. Would the equation be the same if you wanted to plug in any years in it ? without the 30 years ?

OpenStudy (wolf1728):

50,000 *(2^(y/30)) In the formula the y stands for years and you can input any time amount into it. Want to calculate population afetr 3 months? 50,000 *(2^(y/30)) y=.25 and so 50,000 *(2^(.25/30)) = Population = 50,000 *(2^(1/120)) Population = 50,000 * (2^0.0083333333) Population = 50,000 * (2^0.0083333333) Population = 50,000 * 1.0057929411 Population = 50,290

OpenStudy (anonymous):

Okay thank you so much!!!

OpenStudy (wolf1728):

u r welcome